OpenStudy (anonymous):

The function t(x) = x + 6 determines how many cans of soup a food truck needs to stock, where x is the number of shifts the crew is going to work in the truck. The crew uses c(t(x)) to find the amount of money to spend on soup. The function c(x) = 2x + 4. Solve for how much money must be spent when the crew is going to work 4 shifts. a. 24 b. 28 c. 30 d. 34

OpenStudy (anonymous):

a. 24

OpenStudy (anonymous):

how'd you get that?

OpenStudy (anonymous):

you need function composition... substitute t(x) in x of c(x) c(t(x)) = 2(x+6)+4 = 2x+16 then you know that x=4

OpenStudy (anonymous):

did you understand?